# Operatorial formulation of crimo-taxis phenomena in a street

### v1

@@ 4.2 Spatial model
Let us consider a street represented by a one–dimensional region $\mathcal{C}$ divided in $N=50$ adjacent cells, and investigate two scenarios characterized by different spatial distributions of the subgroups. The two situations are described as follows:
@@
there are two areas with a high concentration of ordinary citizens, two suburbs with a small concentration of drug users/dealers, and a very small and dislocated area, with respect to the center of the street, where law enforcement personnel are concentrated. This scenario can be represented by the following initial mean densities:
@@
$n_{2,\alpha}^{0}=0.08\left(\exp(-(\alpha-6)^{2}/44)+\exp(-(\alpha-24)^{2}/26)\right),$
$n_{3,\alpha}^{0}=0.04\exp(-(\alpha-35)^{2}/15),\qquad\alpha\in\mathcal{C};$
@@
there is an area with a high concentration of ordinary citizens, and two small suburban areas occupied by drug users/dealers and law enforcement personnel, respectively. This scenario is similar to the one considered in [1]. We choose the following initial mean densities:
@@
$n_{2,\alpha}^{0}=0.08(\exp(-(\alpha-8)^{2}/42),\qquad\qquad\alpha\in\mathcal{C},$
@@
corresponding to the first and second scenario, respectively. This choice is motivated by the fact that each agent can be thought to have an inertia similar to the initial spatial distribution of the corresponding subgroup.
Finally, we assign the interaction, diffusion and cross-diffusion parameters (as choosen in the model described in [1, 3]), say
@@
Some comments about the numerical simulations are in order. Figures 4 and 5, where first and second scenario, (20) and (21), respectively, are considered, show the time evolution of the mean densities of the three spatially distributed populations, considering both the classical Heisenberg dynamics (Subfigures $(a),(b),(c)$) and the $(\mathcal{H},\rho)$–induced dynamics approach, where the rules (18) and (19) (Subfigures $(d),(e),(f)$ and $(g),(h),(i)$, respectively) are used.
@@
In both scenarios, when classical Heisenberg dynamics is used, we observe an oscillatory trend for all population subgroups, as expected, since no rule is enforced. The results show similarities with Epstein’s model [1], despite the presence of oscillations. In [1], drug users/dealers tend to move toward areas with a higher concentration of ordinary citizens, as these locations offer greater opportunities for committing crimes; law enforcement, in turn, intervenes in response to the presence of drug users/dealers, attempting to reduce their numbers and restore order. This continuous cycle of interaction among criminals, citizens, and law enforcement can generate oscillations in both crime distribution and police intervention areas, similar to the oscillatory behavior observed in Figures 4 and 5. Even in the absence of explicit rules, the feedback among these actors-criminals moving toward citizens, citizens distancing themselves from drug users/dealers, and police reacting can explain the observed dynamics. The phenomenon can be interpreted as a cyclic sequence of phases where crime increases, peaks, and is subsequently countered by police intervention, forming a self-sustaining, dynamically evolving system. It is worth of being remarked that, since the initial inertia parameters have been choosen proportional to the peaks of the initial mean densities of the three subgroups, the tendency of ordinary citizens to remain in their affluent neighborhood is higher with respect to the case of using fixed values for the parameters across the entire domain, as commonly found in the literature [24, 27, 34]. This leads to a significant difference when the classical Heisenberg dynamics is applied: since the inertia parameters assume higher values in areas of higher population densities, the three subgroups show a stronger tendency to remain in their current state at the density peaks. This results in greater resistance to change, leading to the emergence of small stationary pattern formations, although for very short periods of time.
@@
By means of the $(\mathcal{H},\rho)$–induced dynamics approach, we modify the standard Heisenberg dynamics with the rules introduced in Section 3, and the time evolution of the three subgroups accordingly changes (see Figures 4 and 5). By using the rule (18) (see Subfigures $(d),(e),(f)$), involving only modifications of the inertia parameters, the system exhibits a dynamics where oscillations are more controlled, and some patterns remain stable for short periods of time; this behaviour was not possible before. This trend represents a significant first step toward approaching the patterns formation found in [3]. Thus, the obtained outcome can be interpreted as a formation of spatially stationary patterns but dynamically evolving over time, where the population distribution appears stably structured in space but subject to variations and oscillations. This dynamic complexity could represent an intermediate step allowing the formation of more sophisticated and coherent patterns, similar to Turing ones [4], which balance spatial instabilities with time fluctuations. In Subfigures $(g),(h),(i)$ (see Figures 4 and 5), where rules (18) and (19) modify both inertia and diffusion parameters, we note that a trend similar to the case where only the rule (18), acting on inertia parameters, occurs. Anyway, the key difference is that, in this situation, law enforcement personnel are not only attracted towards areas with high criminals concentration but also to areas where ordinary citizens density increases. This fact implies that law enforcement personnel respond not only to zones with high criminal activity but also to densely populated areas, for instance in situations involving public events, demonstrations, or strikes. In other words, their actions are no longer exclusively focused on crime prevention and suppression but also on managing public safety in contexts where a high density of people may occur. This broader and multifaceted approach may reflect a strategy of law enforcement personnel aimed to maintain order and security in high-visibility situations susceptible of rising social risk, thus contributing to a more complex and varied dynamics compared to the model based solely on criminal influence. Additionally, we observe the emergence of stripe-like structures, although some small oscillations persist, which cannot be entirely suppressed as they are intrinsic to the construction of the operatorial model through the quantum-like approach.
@@
Figure 4: First scenario: time evolution and spatial distribution of the three subgroups densities. Ordinary citizens (4, 4 and 4), drug users/dealers (4, 4 and 4), and law enforcement personnel (4, 4 and 4). No-rule (upper), rule (18) (center) and rules (18)–(19) (bottom). The values of initial mean densities and parameters are given in (20), (22) and (24).
@@
Figure 5: Second scenario: time evolution and spatial distribution of the three subgroups densities. Ordinary citizens (5, 5 and 5), drug users/dealers (5, 5 and 5), and law enforcement personnel (5, 5 and 5). No-rule (upper), rule (18) (center) and rules (18)–(19) (bottom). The values of initial mean densities and parameters are given in (21), (23) and (24).
@@ 5 Conclusions
In this paper, we constructed an operatorial model based on fermionic ladder operators to describe the dynamical behavior of three subgroups of a population competing each other and able to diffuse in a one–dimensional spatial region made by a finite number of adjacent cells; the three populations compete locally (in the same cell) and nonlocally (in the adjacent cells). The mean values of the number operators associated to the three agents in a cell are interpreted as a measure of their local density in the cell. The dynamics is ruled by a self–adjoint and time independent quadratic Hamiltonian operator; consequently, adopting the classical Heisenberg viewpoint, the dynamics is ruled by linear ordinary differential equations, and its evolution exhibits a never ending oscillatory trend. To enrich the dynamics, we used the $(\mathcal{H},\rho)$–induced dynamics approach, and suitable choices for the rules are given and motivated. The rules can modify both the inertia and diffusion parameters; however, the cases where the rules act on the inertia parameters only, and on both inertia and the diffusion parameters are investigated. Two scenarios, characterized by different initial spatial distributions of the three subgroups, are considered, and some numerical results, together with their social interpretations, have been presented. The dynamical outcomes show that this model is compatible with the coexistence of the three competing populations in the same environment. The model is susceptible of various extensions and generalizations. For instance, we may include in the Hamiltonian operator effects due to cooperative interactions [34] between the three subgroups of a population, as well as construct a spatial model in a two–dimensional setting, with the aim of investigating if some pattern formations may emerge. Some of these extensions are currently planned and will be the subject of forthcoming papers.
@@ Acknowledgements
This work has been carried out with the patronage of “Gruppo Nazionale per la Fisica Matematica” (GNFM) of “Istituto Nazionale di Alta Matematica” (INdAM). M.G. acknowledges financial support from Finanziamento del Programma Operativo Nazionale (PON) “Ricerca e Innovazione” 2014-2020 a valere sull’Asse IV “Istruzione e ricerca per il recupero”- Azione IV - Dottorati e contratti di ricerca su tematiche dell’innovazione, CUP J45F21001750007. M.G acknowledges financial support from PRIN 2022 “Nonlinear phenomena in low dimensional structures: models, simulations and theoretical aspects”, grant n. 2022TMW2PY. C.F.M. acknowledges financial support from “National Centre for HPC, Big Data and Quantum Computing (HPC)”, Codice progetto CN00000013 – SPOKE 10, CUP D43C22001240001.

### v2

@@ 4.2 Spatial model
Let us consider a street represented by a one–dimensional region $\Omega$ divided in $N=50$ adjacent cells, and investigate two scenarios characterized by different initial spatial distributions of the subgroups. The two situations are described as follows:
@@
there are two areas with a high density of ordinary citizens, two suburbs with a small concentration of drug users/dealers, and a very small and dislocated area, with respect to the center of the street, where law enforcement personnel are concentrated. This scenario can be represented by the following initial mean densities:
@@
$n_{2,\alpha}^{0}=0.08\left(\exp(-(\alpha-6)^{2}/44)+\exp(-(\alpha-24)^{2}/26)\right),$ $\alpha=1,\ldots,50,$
$n_{3,\alpha}^{0}=0.04\exp(-(\alpha-35)^{2}/15);$
@@
there is an area with a high density of ordinary citizens, and two small suburban areas occupied by drug users/dealers and law enforcement personnel, respectively. This scenario is similar to the one considered in [1]. We choose the following initial mean densities:
@@
$n_{2,\alpha}^{0}=0.08(\exp(-(\alpha-8)^{2}/42),\qquad\qquad\alpha=1,\ldots,50,$
@@
corresponding to the first and second scenario, respectively. The choices (22) and (23) describe the fact that each agent can be thought to have an inertia similar to the initial spatial distribution of the corresponding subgroup.
Finally, we assign the interaction, diffusion and cross-diffusion parameters (as chosen in the model described in [1, 3]), say
@@
Some comments about the numerical simulations are in order. Figures 6 and 7, where first and second scenario, (20) and (21), respectively, are considered, show the time evolution of the mean densities of the three spatially distributed populations, considering both the standard Heisenberg view (Subfigures $(a),(b),(c)$) and the $(\mathcal{H},\rho)$–induced dynamics approach, where the rules (18) and (19) (Subfigures $(d),(e),(f)$ and $(g),(h),(i)$, respectively) are used.
@@
In both scenarios, when classical Heisenberg dynamics is used, we observe an oscillatory trend for all subgroups, as expected, since no rule is enforced. The results show similarities with Epstein’s model [1], despite the presence of oscillations. In [1], drug users/dealers tend to move toward areas with a higher concentration of ordinary citizens, as these locations offer greater opportunities for committing crimes; law enforcement, in turn, intervenes in response to the presence of drug users/dealers, attempting to reduce their numbers and restore order. This continuous cycle of interactions among criminals, citizens, and law enforcement can generate oscillations in both crime distribution and police intervention areas, similar to the oscillatory behavior observed in Figures 6 and 7. Even in the absence of explicit rules, the feedback among these agents-criminals moving toward citizens, citizens distancing themselves from drug users/dealers, and police reacting can explain the observed dynamics. The phenomenon can be interpreted as a cyclic sequence of phases where crime increases, peaks, and is subsequently countered by police intervention, forming a self-sustaining, dynamically evolving system. It is worth of being remarked that, since the initial inertia parameters have been choosen proportional to the peaks of the initial mean densities of the three subgroups, the tendency of ordinary citizens to remain in their affluent neighborhood is higher with respect to the case of using fixed values for the parameters across the entire domain, as commonly found in the literature [23, 26, 34]. This leads to a significant difference when the classical Heisenberg dynamics is applied: since the inertia parameters assume higher values in areas of higher population densities, the three subgroups show a stronger tendency to remain in their current state at the density peaks. This results in greater resistance to change, leading to the emergence of small stationary pattern formations, although for very short periods of time.
@@
By means of the $(\mathcal{H},\rho)$–induced dynamics approach, the standard Heisenberg dynamics is modified through the rules introduced in Section 3, and the time evolution of the three subgroups accordingly changes (see Figures 6 and 7). By using the rule (18) (see Subfigures $(d),(e),(f)$), involving only modifications of the inertia parameters, the system exhibits a dynamics where oscillations are more controlled, and some patterns remain stable for short periods of time; this behaviour was not possible before. This trend represents a significant first step toward approaching the patterns formation found in [3]. Thus, the obtained outcome can be interpreted as a formation of spatially stationary patterns but dynamically evolving over time, where the population distribution appears stably structured in space but subject to variations and oscillations. This dynamic complexity could represent an intermediate step allowing the formation of more sophisticated and coherent patterns, similar to Turing ones [4], which balance spatial instabilities with time fluctuations. In Subfigures $(g),(h),(i)$ (see Figures 6 and 7), where rules (18) and (19) modify both inertia and diffusion parameters, we note that a trend similar to the case where only the rule (18), acting on inertia parameters, occurs. Anyway, the key difference is that, in this situation, law enforcement personnel are not only attracted towards areas with high criminals concentration but also to areas where ordinary citizens density increases. This because law enforcement personnel can be attracted not only toward zones with high criminal activity but also to densely populated areas, for instance in situations involving public events, such as demonstrations, or strikes. In other words, their actions are no longer exclusively focused on crime prevention and suppression but also on managing public safety in contexts where a high density of people may occur. This broader and multifaceted approach may reflect a strategy of law enforcement personnel aimed to maintain order and security in high-visibility situations susceptible of rising social risk, thus contributing to a more complex and varied dynamics compared to the model based solely on criminal influence. Additionally, we observe the emergence of stripe-like structures, although some small oscillations persist, which cannot be entirely suppressed as they are intrinsic to the construction of the operatorial model through the quantum-like approach.
@@
Figure 6: First scenario: time evolution and spatial distribution of the three subgroup densities. Ordinary citizens (6, 6 and 6), drug users/dealers (6, 6 and 6), and law enforcement personnel (6, 6 and 6). No-rule (upper), rule (18) (center) and rules (18)–(19) (bottom). The values of initial mean densities and parameters are given in (20), (22) and (24).
@@
Figure 7: Second scenario: time evolution and spatial distribution of the three subgroup densities. Ordinary citizens (7, 7 and 7), drug users/dealers (7, 7 and 7), and law enforcement personnel (7, 7 and 7). No-rule (upper), rule (18) (center) and rules (18)–(19) (bottom). The values of initial mean densities and parameters are given in (21), (23) and (24).
@@ 5 Conclusions
In this paper, we implemented an operatorial version of a model for crime phenomena, originally introduced by Epstein [1] and further investigated in [3]. The population is partitioned in three subgroups that are modeled by fermionic ladder operators whose time evolution is assumed to be ruled by a Hermitian time-independent quadratic Hamiltonian operator. The three subgroups exhibit either a local or a nonlocal competition, and are able to spread in a one–dimensional spatial region. The mean values of the number operators corresponding to the three subgroups in a cell can be interpreted as a measure of their local density in the cell. In order to have a more realistic dynamics, we used the $(\mathcal{H},\rho)$–induced dynamics approach, and suitable choices for the rules are given and motivated. The rules can modify the inertia and diffusion parameters, according to the evolution of the system. Two scenarios, characterized by different initial spatial distributions of the three subgroups, are considered, and some numerical results, together with their social interpretation, have been presented. Several extensions and generalizations of this model can be considered: effects due to cooperative interactions among the three subgroups can be incorporated in the Hamiltonian operator [34], as well as a spatial model in a two–dimensional setting, with the aim of investigating if some stable patterns may emerge, can be considered. Work is in progress in these directions.
@@ Acknowledgements
This work has been carried out with the patronage of “Gruppo Nazionale per la Fisica Matematica” (GNFM) of “Istituto Nazionale di Alta Matematica” (INdAM). M.G. acknowledges financial support from Finanziamento del Programma Operativo Nazionale (PON) “Ricerca e Innovazione” 2014-2020 a valere sull’Asse IV “Istruzione e ricerca per il recupero”- Azione IV - Dottorati e contratti di ricerca su tematiche dell’innovazione, CUP J45F21001750007. M.G acknowledges financial support from PRIN 2022 “Nonlinear phenomena in low dimensional structures: models, simulations and theoretical aspects”, grant n. 2022TMW2PY. G.I. acknowledges financial support from the project “Reclutamento di Ricercatori di tipo A ed Assegnisti di Ricerca, Università degli Studi di Messina (Regione Sicilia)”, CUP G41I22000690001. C.F.M. acknowledges financial support from “National Centre for HPC, Big Data and Quantum Computing (HPC)”, Codice progetto CN00000013 – SPOKE 10, CUP D43C22001240001.